OpenStudy (.whitedragon.):

NEED NOW !!What is the value of y so that the ordered pair (2, y) is a solution to the equation 2x+y=-2 ?

OpenStudy (.whitedragon.):

pls hlp

OpenStudy (.whitedragon.):

A. –6 B. –2 C. -1/2 D. 0

OpenStudy (anonymous):

Substitute 2 in for x. 2(2)+y=-2 4+y=-2 now do inverse operation to get y by itself. In this case, you will subtract 4 to both sides. y=-6 Viola! Check by substituting both numbers into its rightful place. 2(2)+(-6)=-2 -2=-2 Hope that helps! Ask me more if you need help:)